A sample of argon gas at a pressure of 0.871 atm and a temperature of 26.7 °C, occupies a volume of 809 mL. If the gas is compressed at constant temperature until its pressure is 1.28 atm, the volume of the gas sample will be

________ mL.
1
Expert's answer
2019-03-13T07:00:52-0400

P1V1 = P2V2

V2 = P1V1/P2

V2 =  0.871*809/1.28 = 550.5 (ml)